Math 215 Section 2 Fall 95 Homework Assignments

Assignment Date Reading Problems  Due Date
9/18  1.3  2, 4, 8, 10, 11, 13  9/27
9/20  2.7  2, 4, 6, 8, 9  9/27
9/22  3.1  5-8, 15, 16, 18, 19, 21-24  9/27
Computer HW 4 and 5  9/27

Notes: Part c of 1.3 #13 is not optional.

9/25  3.2  6, 7, 13, 15, 21-22  10/4
9/27  3.3  2-10 evens, 14, 16, 18, 20  10/4
9/29  3.5  2, 8, 10, 17, 26  10/4
Computer Projects 3 and 5  10/4

Notes: Problem 14 of Section 3.3 should read “has only the trivial solution.”

General Note: Ranges of numbers like 5-8 or 21-22 mean that the problems go together and will count as one problem in grading. When the book mentions MATLAB, you can use Linalg instead.